OpenStudy (anonymous):

A line has a slope of -5 and contains the point (-2, 4). Which equations represent the line? Choose all answers that are correct

OpenStudy (anonymous):

A. -5x + y = -6 B. 5x + y = -6 C. (y - 4) = -5(x + 2) D. (y + 4) = -5(x - 2)

OpenStudy (anonymous):

@haleyelizabeth2017

geerky42 (geerky42):

Recalling point-slope form: \(y-y_1 = m(x-x_1)\), where \(m\) is slope and \((x_1,y_1)\) is point line goes through.

geerky42 (geerky42):

You have \(m = -5\), \(x_1 = -2\) and \(y_1 = 4\)